$SPAD/src/input uniseg.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool uniseg.output
)set message test on
)set message auto off
)clear all
--S 1 of 9
pints  := 1..
--R
--R
--R   (1)  1..
--R                                      Type: UniversalSegment(PositiveInteger)
--E 1
--S 2 of 9
nevens := (0..) by -2
--R
--R
--R   (2)  0.. by - 2
--R                                   Type: UniversalSegment(NonNegativeInteger)
--E 2
--S 3 of 9
useg: UniversalSegment(Integer) := 3..10
--R
--R
--R   (3)  3..10
--R                                              Type: UniversalSegment(Integer)
--E 3
--S 4 of 9
hasHi pints
--R
--R
--R   (4)  false
--R                                                                Type: Boolean
--E 4
--S 5 of 9
hasHi nevens
--R

--R
--R   (5)  false
--R                                                                Type: Boolean
--E 5
--S 6 of 9
hasHi useg
--R
--R
--R   (6)  true
--R                                                                Type: Boolean
--E 6
--S 7 of 9
expand pints
--R
--R
--R   (7)  [1,2,3,4,5,6,7,8,9,10,...]
--R                                                        Type: Stream(Integer)
--E 7
--S 8 of 9
expand nevens
--R
--R
--R   (8)  [0,- 2,- 4,- 6,- 8,- 10,- 12,- 14,- 16,- 18,...]
--R                                                        Type: Stream(Integer)
--E 8
--S 9 of 9
expand [1, 3, 10..15, 100..]
--R
--R
--R   (9)  [1,3,10,11,12,13,14,15,100,101,...]
--R                                                        Type: Stream(Integer)
--E 9
)spool
)lisp (bye)

———-

References^{[1] nothing}

5